OpenStudy (liizzyliizz):

find dy/dx by implicit differentiation for x^1/2 + y^1/2 = 16 I kind of understand the guidelines for it, but at the same time I'm ehh about it.

OpenStudy (anonymous):

ok so since they want you to use implicit differentiation, you don't need to solve for "y" first...

OpenStudy (anonymous):

basically you just do this: 1/2x^-1/2 + 1/2y^-1/2dy/dx = 0

OpenStudy (liizzyliizz):

i worked it out and i checked the answer and it says that it should be \[-\sqrt{y/x} \] im not sure how that happened.

OpenStudy (anonymous):

you have to solve for y prime by making the dy/dx stand for y prime.

OpenStudy (anonymous):

all you do is take the derivatives of each piece like you normally would, but when you take the derivative of "y", you need to remember the dy/dx part...because you're eventually solving for dy/dx

OpenStudy (anonymous):

do you understand what i did so far?

OpenStudy (anonymous):

the derivative of x^1/2 is 1/2x^-1/2

OpenStudy (liizzyliizz):

yeah that i understand

OpenStudy (liizzyliizz):

the 1/2 in the front is cancelled out? that just dissapeared. thats what i wanted to know.

OpenStudy (liizzyliizz):

Im saying that for a later step, umm im going through the work for the answer online and i didnt understand what happened there.

OpenStudy (amistre64):

\[D[x^{1/2} + y^{1/2} = 16]\] \[D[x^{1/2}] +D[y^{1/2}] = D[16]\] \[\frac{1}{2x^{1/2}}x' +\frac{1}{2y^{1/2}}y' = 0\] \[\frac{1}{2y^{1/2}}y' = -\frac{1}{2x^{1/2}}x'\] \[y' = -\frac{2y^{1/2}}{2x^{1/2}}x'\] \[y' = -\frac{y^{1/2}}{x^{1/2}}x';\ \text{and x' = 1 with respect to x}\] \[y' = -\frac{y^{1/2}}{x^{1/2}}\]

OpenStudy (anonymous):

my internet sucks right now :(

OpenStudy (liizzyliizz):

it's ok im still trying to understand amistre, i feel like i understand this , like as far as the guidelines that my textbook explain. Now even better with amistre's explanation and everyone elses. But if I continue doing this, i'm still going to make some sort of mistake... -.-

OpenStudy (amistre64):

implicits are nothin special, or any different from explicits

OpenStudy (amistre64):

tell me, what is the derivative of: 3x^2 ?

OpenStudy (liizzyliizz):

6x

OpenStudy (amistre64):

close, but no. 6x x' i never said the derivative with respect to x. and what you are used to doing is throwing this precious little derived bit away since it equals 1; with respect to itself

OpenStudy (amistre64):

what is the derivative of: 3y^2 ? ; or 3r^2, or 3m^2?

OpenStudy (amistre64):

the rules dont change and never have; you are just used to throwing out the derived bit by habit instead of sorting it out in the end

OpenStudy (liizzyliizz):

so it would be 6y y' ?

OpenStudy (amistre64):

yes

OpenStudy (liizzyliizz):

hmm maybe because i just stated doing this and my teacher didnt explain that, but she never told me to do that and my text book doesnt show that. (atleast so far... )

OpenStudy (amistre64):

it wont show that either ....

OpenStudy (liizzyliizz):

oh so thats just something i should know? hmmm

OpenStudy (amistre64):

its something you should be taught :) and once taught, all the rest of the enigma tends to go away

OpenStudy (amistre64):

the product rule is still the product rule right? [xy]' = x'y+xy'

OpenStudy (liizzyliizz):

yes lol

OpenStudy (amistre64):

if this is with respect yo y; y'=1; if with respect to x, x'=1; if respect to time; neither equals 1

OpenStudy (amistre64):

these derived bit are the workings of the chain rule; always poping out little derived bits

OpenStudy (amistre64):

keep them and sort thru them in the end and you will never go wrong

OpenStudy (liizzyliizz):

So I would use the chain rule for lets say x^2y+y^2x =-2 I've been kind of stuck on how to approach this one

OpenStudy (amistre64):

the chain rule is at work; but not the full blown method that youre used to it being here

OpenStudy (amistre64):

i see product, and power rules

OpenStudy (amistre64):

D[x^2 y + y^2 x = -2] D[x^2 y] + D[y^2 x] = D[-2] -------------------------------------- D[x^2 y] = x'^2 y+x^2 y' = 2x x' y + x^2 y' D[y^2 x] = y^2' x+y^2 x' = 2y y' x+ y^2 x' D[-2] = 0 ---------------------------- 2x x' y + x^2 y' + 2y y' x+ y^2 x' = 0 and since x'=1 with respect to x we can toss them 2xy + x^2 y' + 2y y' x+ y^2 = 0 and solve for y'

OpenStudy (amistre64):

x^2 y' + 2xy y' = -2xy - y^2 y' (x^2 + 2xy) = -2xy - y^2 y' = (-2xy - y^2) ---------- (x^2 + 2xy)

OpenStudy (amistre64):

or neatly: \[D[x^2 y + y^2 x = -2]\] \[D[x^2 y] + D[y^2 x] = D[-2]\] -------------------------------------- \[D[x^2 y] = x'^2 y+x^2 y' = 2x x' y + x^2 y'\] \[ D[y^2 x] = y'^2 x+ y^2 x' = 2y y' x+ y^2 x' \] \[D[-2] = 0\] ---------------------------- \[2x x' y + x^2 y' + 2y y' x+ y^2 x' = 0\] and sort

OpenStudy (liizzyliizz):

Ohh boy, and to think i was once good at math, this has made me think twice about it. So for some reason the whole x' y' is totally throwing me off.

OpenStudy (amistre64):

they are simply place holders in a sense

OpenStudy (amistre64):

you can reformat them as \(y'=\cfrac{dy}{d*}\) and \(x'=\cfrac{dx}{d*}\) and then fill in the * part afterwards really

OpenStudy (amistre64):

if we say we want this with respect to x then we get: \(y'=\cfrac{dy}{dx}\) and \(x'=\cfrac{dx}{dx}=1\)

OpenStudy (amistre64):

the best thing to do is to pick a method that works best for you and own it :)

OpenStudy (liizzyliizz):

maybe thats whats throwing me off, i use dy/dx and dx/dx which is 1 instead of y' and it just looks weird to me. because based on the work u showed me i think i figured how to do it. but im not sure

OpenStudy (liizzyliizz):

woah not because i meant but * lol